You just sold short 1,000 shares of Wetscope, Inc., a fledgling software firm, at $58 per share. You cover your short when the price hits $48.50 per share one year later. If the company paid $0.63 per share in dividends over this period, what is your percentage rate of return on the investment? Assume an initial margin of 55 percent.

Rate of return?
Answer #1

The initial investment we keep is according to margin which is 55%

initial investment=(no of shares short* price)*55%

=1000*58*55%=31900

Final return we got after one year=(no of shares*(short price))-(dividend income+(no of shares*final price))

=(1000*58)-((1000*0.63)+(1000*48.5))

=8870

Return=(final/initial)

=(8870/31900)

=27.81%